A saline solution contains 0.90 g NaCl in exactly 100. mL of solution. What is the molarity of the solution ?
1
Expert's answer
2019-04-01T07:47:34-0400

"C_m= \frac{m}{M \cdot V}"

Cm - molarity or molar concentration of the solution

m - mass of the solute, g

M - molar mass of the solute, g/mol

V - volume of the solution, L


"M(NaCl)=1 \cdot 22.99 + 1 \cdot 35.453 = 58.443 \; g/mol"

"C_m= \frac{0.9}{58.443 \cdot 0.1} \approx 0.15 \; M"


Answer: 0.15 M.